find the area enclosed by the curve r²-4cosø


1
Expert's answer
2021-06-07T15:53:44-0400

"\displaystyle
r^2 = 4\cos{\phi}\\
\begin{aligned}
A &= 2 \times \frac{1}{2} \int_{-\frac{\pi}{2}}^{\frac{\pi}{2}} 4\cos{\phi} \,\,\mathrm{d}\phi 
\\&= 4 \sin{\phi}\vert_{-\frac{\pi}{2}}^{\frac{\pi}{2}}
\\&= 8
\end{aligned}"